solve for x tan^-1(x/2) + tan^-1(x/3) = pi/4 ,0

solve for x
tan^-1(x/2) + tan^-1(x/3) = pi/4 ,0< x< root 6

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
take tan both side

tan(tan^-1(x/2) + tan^-1(x/3))=tan(pi/4)
(x/2+x/3)/(1-x^2/6)=1
x/2+x/3=1-x^2/6

now solve for x

sher mohammad